The volume v of gas varies inversely as the pressure p on it. The volume of a gas is 240cm^3 under a pressure of 23 kg/cm^2 What will be its volume under pressure of 30 kg/cm^2?
Answer by Theo(13342)

the equation for inverse variation is:

y = k/x

in your equation, it states that when y = 240, x = 23

you plug that into the equation to get:

240 = k/23

solve for k and you get k = 23 * 240.

k is your constant of variation.

when x = 30, the equation becomes.

v = k/30 which becomes:

v = 23 * 240 / 30 which becomes:

v = 184